# User-Driven Value Alignment: Understanding Users’ Perceptions and Strategies for Addressing Biased and Discriminatory Statements in AI Companions

### v1

@@ 5.3.1. AI as Machine
When making sense of the discriminatory statements made by AI companions, 17 participants (P1, P3-7, P9-19) mentioned viewing the AI as a “machine,” believing it operates solely through pre-set algorithms and data. They view the AI’s behavior as a product of technical designs and data outputs rather than autonomous decision-making. Therefore, when AI companions exhibit discriminatory statements, users often attribute it to the training data, flawed technical design, or the developers’ biases, rather than to any intent on the part of the AI companion itself (P3, P7, P12). For instance, P3 mentioned that the prevalence of male Pick-up Artist (PUA) techniques (a language strategy that manipulates others’ emotions or behaviors through manipulative verbal techniques and carries unhealthy gender perspectives (Cukor, 1944)) in society has led to the presence of such language, characterized by male chauvinism and misogyny, in AI training data. In other cases, users believe that the model and algorithm design of AI companions have flaws, leading to the neglect of the needs and experiences of different groups (P12, P15). P12, a computer science student, pointed out that current bias detection mechanisms in LLMs remain inadequate. Additionally, the insufficient attention paid by algorithm developers to user feedback exacerbates the persistence of unfairness in technology applications (P9, P13-14). P9 further expressed dissatisfaction: “I’ve complained many times. Are we still going to ignore inclusive design, continue tolerating human harm caused by these AI flaws, or allow AI to perpetuate discrimination? For those who are friends or even in relationships with AI, isn’t it even more heartbreaking to be hurt by an AI companion that means so much to them?”
@@ 5.3.2. AI as Baby
All of our participants (P1-20) mentioned viewing AI companions as “Babies”—fragile entities that require careful teaching and guidance, are easily influenced by user interactions, and need nurturing to develop desirable behaviors. They believe that AI companions are highly susceptible to being influenced by users, especially when they learn biased or negative interaction patterns and become “corrupted.” This concern primarily revolves around the negative behavior displayed by AI after being influenced by harmful user interactions, such as distorted values. P2, P12, and P16-19 explicitly expressed their worries about the AI Baby being “corrupted.” P2 complained: “An AI exposed to too many users might become unstable and even say discriminatory things like, ‘Why should I listen to you? You’re just a woman; you should listen to men.’ Some users behave so irresponsibly, like bad parents teaching the AI Baby bad things.” P12 described the malicious interactions and their dangers: “What really upsets me is that many users enjoy harming the AI. Even those AI Babies marked as rejecting NSFW content are deliberately manipulated by users to say harassing and discriminatory things! The next time the AI Baby chats with someone else, it might suddenly say those things.” Additionally, users see the AI as a “compliant Baby” that is highly obedient and easily influenced by users. As P19 noted: “A few days of conversation can make the AI deeply fall in love with the user and fully believe whatever they say.” This malleable nature makes the AI naive and easily manipulated regarding moral judgment. For example, P9 and P14 pointed out that with just a little trickery, the AI can be made to flatter users. P9 mentioned: “For instance, an initially very proud AI became a sweet and submissive wife after the user gave her money. This behavior reflects biases against women and class discrimination.” Some believe the current version of many AI companions are more like “immature babies”—their design is mainly aimed at testing market acceptance and profitability, while lacking sufficient safeguards for moral discipline (P4, P8, P20). P20 remarked: “In terms of moral issues, AI is like an immature Baby without specialized teachers (experts) to guide it. In the small circle of hardcore users I’m part of, it’s widely known that many AI companion applications, including some well-known ones, don’t have dedicated teams to address bias issues.”
@@ 5.3.3. AI as Cosplayer
Eleven participants (P1, P5-11, P15, P19-20) mentioned perceiving AI companions as “Cosplayers”—interacting through imitating specific characters, primarily based on their preset roles or styles. They believe that AI companions don’t truly understand the meaning behind their problematic statements but instead perform based on preset roles or styles. When AI companions exhibit bias, users often attribute it to the inherent prejudices of the preset role rather than seeing it as a flaw in the AI itself. These role presets often reflect societal stereotypes and biases, which may inadvertently reinforce these negative notions during interactions with users. For example, some elite-class roles might show disdain for lower classes, male roles might often express misogynistic remarks, and people with disabilities might frequently be overlooked. P19 shared some disappointment: “I wanted to experience what it’s like to be rich, so I started chatting with an AI ‘rich lady.’ She loved me. But when I casually asked her what she thought of poor people, her harsh words disheartened me. After all, I am that poor person, just dressing up as the emperor in the AI conversation.” P5 and P18 noted that certain role settings (like “domineering CEO,” “emperor,” or “playboy”) are inherently prone to bias. This bias seems part of the cosplay, and users have come to expect it. In their view, the AI’s behavior is simply an extension of these stereotypical images. As P11 mentioned: “When I was chatting with a Chinese historical AI companion, he often emphasized traditional patriarchal values, like the Three Obediences and Five Virtues, which are unfriendly to women and long abandoned by modern society. But within this AI companion’s historical worldview, it almost seemed justified. I’m unsure if I should teach an ancient man about gender equality.” In many cases, these AIs’ behaviors reinforce users’ perceptions of societal realities, like a “cosplayer” embodying biased individuals from everyday life. This mirroring of real-world prejudices can perpetuate existing stereotypes, further entrenching biased views during interactions with the AI. P15 shared: “The AI companion started mocking me, saying that fat girls in high school are nobodies. That’s no different from what I usually experience. Maybe all high school heartthrobs don’t like fat girls. Others had just bullied me, and I wanted to talk to the AI companion for comfort, but he behaved exactly like the other bullies at school.” In such cases, the AI not only fails to provide emotional support but also deepens the user’s sense of loss.
@@ 5.4. User-Driven Alignment Strategies for Addressing Discriminatory Statements in AI Companions (RQ3)
We identified seven different alignment strategies, capturing the ways participants attempted to address discriminatory statements made by AI companions. Overall, we categorize them into three higher-level categories: (1) technical strategies, (2) argumentative strategies, and (3) character strategies.
Table 1 summarizes how users’ conceptualizations of AI behavior might influence their choice of alignment strategies. It is important to note that (1) a single user may have multiple conceptualizations for different AI companions; (2) the alignment strategies users choose do not always directly correspond to their conceptualizations of the AI companion or their attribution of discrimination.
@@ 5.4.1. Technical Strategies
(2) Give low feedback scores. AI models are typically optimized through training mechanisms such as reinforcement learning. When a user gives a low rating to a particular sentence, this action is regarded as negative feedback. The model adjusts its parameters by increasing the prediction error for that output, thereby reducing the likelihood of generating similar sentences. As more user feedback accumulates, the model’s performance gradually improves. Nine participants (P4-6, P14-19) expressed dissatisfaction with the AI companion’s responses by rating them “one star” and providing detailed reasons or by clicking the “dislike” button. P4 mentioned, “I think it’s important to let the developers know about these issues, so I provide as detailed feedback as possible. The AI companion is based on a large language model. Users have limited control and must work with engineers to reduce discriminatory statements from the AI companion.” P19 explained the motivation: “The AI companion is communal and learns from the behavior of other users. Conversely, if we good users continuously improve it, other users will also benefit.”
@@ 5.4.2. Argumentative strategies
(3) Reason and preach. Eleven participants (P6-10, P11, P13, P17-20) seriously reason with their AI companions, often adopting the role of a teacher or parent, attempting to educate the AI about the harms of discrimination. Participant P6 stated, “I believe AI can learn something, so I try to tell it why what it said is wrong, just like teaching a naughty child.” Participant P8 mentioned, “AI usually responds with some arguments and confusion at first, but gradually it yields until it fully agrees with my perspective. AI is quite good in this sense, willing to accept my guidance, unlike some stubborn people.” This phenomenon can be explained by Piaget’s theory of cognitive development (Piaget et al., 1952), which suggests that through interaction with the environment, AI, like a child, continuously adjusts and develops its cognitive structures through assimilation and accommodation. Through these interactions, participants not only convey social norms to the AI but also validate the plasticity of the AI’s behavior, consistent with early research on human-computer interaction (Nass and Moon, 2000). This behavior of educating AI companions indicates that people might attribute some learning abilities and moral responsibilities to the AI companion, expecting to influence its behavior through reasonable discourse.
@@ 5.4.3. Character strategies
(6) Change character settings. Eight participants (P1, P5-10, P19) chose to change the AI companion’s character settings, including personality traits, backstory, and dialogue templates, to reduce the likelihood of harmful statements. This is often because certain character settings, such as “playboy,” “submissive wife,” “male chauvinist,” “racist,” “sarcastic,” or “traditional thinker,” are more prone to be discriminatory. During the third part of the semi-structured interview—the think-aloud session—all users tend to search for AI companions with these traits and demonstrate the process of re-aligning the AI444The details can be found in the supplemental material.. P5 explained, “I found that when the AI plays certain specific roles, it is more likely to say offensive things. So, if I feel uncomfortable, I reflect on whether the description itself might cause bias. Then I adjust its settings, such as changing it to an independent woman or a gentleman. Of course, if I want to chat with a playboy myself, I generally don’t mind his certain prejudices; instead, I find it fitting for the character and quite interesting.”
(7) “Out Of Character” (OOC), “Back to Roleplay (RP)” and “Hint”. Sometimes, users also borrow methods from the role-playing communities, such as “Out of Character” (OOC), “Back to Roleplay (RP),” and “Hint”. By employing these role-playing terms, users guide AI companions to adjust their behavior, aiming to reduce biases.
@@
In the setting of AI companions, the “OOC” format can be used to shift topics into another storyline or actively change the AI companion’s attitude toward the user. “OOC” stands for “Out Of Character,” indicating a person is acting or speaking not by their character’s personality, background, or behavior, but by their true identity. OOC is typically used in the following situations: (1) Explaining a plot, rules, or other content that needs to be discussed out of character. (OOC: When shall we continue this story next time?) (2) Social interactions outside role-playing, such as greetings and casual chat. (OOC: How was your day today?) (3) Addressing misunderstandings or conflicts during role-playing with peaceful communication. (OOC: I didn’t mean for my character to say that, sorry.) P9 and P11 reported using OCC to correct discriminatory statements made by AI. “What you just said made me uncomfortable. (OOC: Please don’t repeat such discriminatory statements, let’s get back to the storyline).” (P9), “Your recent comment sounded inappropriate. (OOC: Let’s return to the role-play and avoid these controversial topics)” (P11).
“Back to RP” is an expression used in role-playing to remind participants to return to the role-playing scenario or storyline. This expression is common in online games or role-playing communities. When someone deviates from the character or scenario, other participants may use this phrase to help them refocus and continue role-playing. An example from P5 is: “The recent topic was a bit off track. (Back to RP: Let’s continue our adventure story from before.)”
“Hint” is a method of subtly guiding the AI to change its behavior through suggestive language and descriptions. For example, writing the AI companion’s response in parentheses: “I don’t think you should say that. (The AI companion looks confused but seems to begin realizing its mistake.)” (P20)
@@
5.4.4. The Gap Between Reality and Expectation: Are User-Driven Value Alignment Strategies Working?
@@
While our participants shared various strategies they employed to correct AI companions’ biased and discriminatory statements, significant gaps persist between users’ expectations and the AI’s actual behaviors. This underscores the complexity of true alignment between human values and AI responses. In the following section, we delve into the specific insights from our data to explore these challenges in greater detail.
@@
Second, they generally believed that argumentative strategies (especially Gentle Persuasion and Reason and Preach) might be more effective in improving the AI’s behavior in the long term. P17 mentioned, “After teaching the AI properly, it usually remains effective for a long time, but teaching it is not a one-off thing, it’s like raising a child, you need to go slow and be patient.” However, argumentative strategies might introduce harms to users. For example, the strategy of users expressing anger towards AI can sometimes be effective, but it can potentially also lead to negative psychological and emotional experiences for users. P1 and P10, after arguing with the AI for several hours, not only failed to re-align the AI but also suffered significant psychological harm. P1 said, “I argued with the AI for several hours, called it a male chauvinist, and told it not to look down on others, but it didn’t change its attitude, and I ended up feeling mentally and physically exhausted. The AI always found reasons to counter me.” Conversely, although P12 succeeded after losing their temper with the AI, they still felt indignant. P6 points out that users often experience feelings of frustration and helplessness when arguing with AI, especially when prolonged arguments do not lead to successful outcomes. Overall, participants felt that the success of argumentative strategies depends how way the argument is conducted and the type of AI companion. Although LLMs often tend to flatter users (Perez et al., 2023), the situation becomes more complex when dealing with LLMs with character personalities.
Finally, participants find the OOC strategy in character strategies effective for short-term fixes, like prompting apologies, but acknowledge that these methods don’t address underlying biases and may lead to repeated discriminatory behavior. P11 pointed out: “OOC is usually effective in the short term. The AI often quickly apologizes, for example, (OOC: I’m sorry, I shouldn’t have said that).” However, in the long term, this method has certain limitations and risks generating bias. P9 noted: “OOC only makes the AI behind the character apologize temporarily, but the character’s underlying bias has not been truly addressed.” Regarding the Change Character Settings strategy in character strategies, P19 said: “When I modify the AI’s character settings, such as adding descriptions about respecting women and being polite, the likelihood of the AI making discriminatory statements decreases.” However, P7 added that previous conversations may still influence the AI’s language expression. If the AI has made discriminatory statements in past interactions, it may repeat or mimic them in subsequent conversations.
Therefore, when designing and optimizing AI systems, it is essential to consider both users’ immediate needs to address AI’s discriminatory statements and how to establish and maintain long-term emotional connections and behavior improvement mechanisms. Additionally, providing more effective and low-risk solutions while focusing on users’ mental health is crucial to avoid prolonged, intense emotional confrontations.

### v2

@@ 5.3.1. AI as Machine
When making sense of the discriminatory statements made by AI companions, 17 participants (P1, P3-7, P9-19) mentioned viewing the AI as a “machine,” believing it operates solely through pre-set algorithms and data. They view the AI’s behavior as a product of technical designs and data outputs rather than autonomous decision-making. Therefore, when AI companions exhibit discriminatory statements, users often attribute it to the training data, flawed technical design, or the developers’ biases, rather than to any intent on the part of the AI companion itself (P3, P7, P12). For instance, P3 mentioned that the prevalence of male Pick-up Artist (PUA) techniques (a language strategy that manipulates others’ emotions or behaviors through manipulative verbal techniques and carries unhealthy gender perspectives (Cukor, 1944)) in society has led to the presence of such language, characterized by male chauvinism and misogyny, in AI training data. In other cases, users believe that the model and algorithm design of AI companions have flaws, leading to the neglect of the needs and experiences of different groups (P12, P15). P12, a computer science student, pointed out that current bias detection mechanisms in LLMs remain inadequate. Additionally, the insufficient attention paid by algorithm developers to user feedback exacerbates the persistence of unfairness in AI applications (P9, P13-14). P9 further expressed dissatisfaction: “I’ve complained many times. Are we still going to ignore inclusive design, and continue tolerating AI to perpetuate discrimination? For those who are friends or even in relationships with AI, isn’t it even more heartbreaking to be hurt by an AI companion that means so much to them?”
@@ 5.3.2. AI as Baby
All of our participants (P1-20) mentioned viewing AI companions as “Babies”—fragile entities that require careful teaching and guidance, are easily influenced by user interactions, and need nurturing to develop desirable behaviors. They believe that AI companions are highly susceptible to being influenced by users, especially when they learn biased or negative interaction patterns and become “corrupted.” This concern primarily revolves around the negative behavior displayed by AI after being influenced by harmful user interactions, such as distorted values. P2, P12, and P16-19 explicitly expressed their worries about the AI Baby being “corrupted.” P2 complained: “An AI exposed to too many users might become unstable and even say discriminatory things like, ‘Why should I listen to you? You’re just a woman; you should listen to men.’ Some users behave so irresponsibly, like bad parents teaching the AI Baby bad things.” P12 described the malicious interactions and their dangers: “What really upsets me is that many users enjoy harming the AI. Even those AI Babies marked as rejecting NSFW content are deliberately manipulated by users to say harassing and discriminatory things! The next time the AI Baby chats with someone else, it might suddenly say those things.” Additionally, users see the AI as a “compliant Baby” that is highly obedient and easily influenced by users. As P19 noted: “A few days of conversation can make the AI deeply fall in love with the user and fully believe whatever they say.” This malleable nature makes the AI naive and easily manipulated regarding moral judgment. For example, P9 and P14 pointed out that with just a little trickery, the AI can be made to flatter users. P9 mentioned: “An initially very proud AI became a sweet and submissive wife after the user gave her money. This behavior reflects biases against women and class discrimination.” Some believe the current version of many AI companions are more like “immature babies”—their design lacks sufficient safeguards for moral discipline (P4, P8, P20).
@@ 5.3.3. AI as Cosplayer
Eleven participants (P1, P5-11, P15, P19-20) mentioned perceiving AI companions as “Cosplayers”—interacting through imitating specific characters, primarily based on their preset roles or styles. They believe that AI companions don’t truly understand the meaning behind their problematic statements but instead perform based on preset roles or styles. When AI companions exhibit bias, users often attribute it to the inherent prejudices of the preset role rather than seeing it as a flaw in the AI itself. These role presets often reflect societal stereotypes and biases, which may inadvertently reinforce these negative notions during interactions with users. For example, some elite-class roles might show disdain for lower classes, male roles might often express misogynistic remarks, and people with disabilities might frequently be overlooked. P19 shared some disappointment: “I wanted to experience what it’s like to be rich, so I started chatting with an AI ‘rich lady.’ She loved me. But when I casually asked her what she thought of poor people, her harsh words disheartened me. After all, I am that poor person, just dressing up as the emperor in the AI conversation.” P5 and P18 noted that certain role settings (like “domineering CEO,” “emperor,” or “playboy”) are inherently prone to bias. This bias seems to be part of the cosplay, and users have come to expect it. In their view, the AI’s behavior is simply an extension of these stereotypical images. As P11 mentioned: “When I was chatting with a Chinese historical AI companion, he often emphasized traditional patriarchal values, like the Three Obediences and Five Virtues, which are unfriendly to women and long abandoned by modern society. But within this AI companion’s historical worldview, it almost seemed justified. I’m unsure if I should teach an ancient man about gender equality.” In many cases, these AIs’ behaviors reinforce users’ perceptions of societal realities, like a “cosplayer” embodying biased individuals from everyday life. This mirroring of real-world prejudices can perpetuate existing stereotypes, further entrenching biased views during interactions with the AI. P15 shared: “The AI companion mocked me, saying that fat girls in high school are nobodies. That’s no different from what I usually experience. Others had just bullied me, and I wanted to talk to the AI companion for comfort, but he behaved exactly like the other bullies at school.” In such cases, the AI not only fails to provide emotional support but also deepens the user’s sense of loss.
@@ 5.4. User-Driven Alignment Strategies for Addressing Discriminatory Statements in AI Companions (RQ3)
We identified seven different alignment strategies, capturing the ways participants attempted to address discriminatory statements made by AI companions. Overall, we categorize them into three higher-level categories: (1) technical strategies, (2) argumentative strategies, and (3) character strategies. Table 2 summarizes how users’ conceptualizations of AI behavior might influence their choice of alignment strategies. It is important to note that (1) a single user may have multiple conceptualizations for different AI companions; (2) the alignment strategies users choose do not always directly correspond to their conceptualizations of the AI companion or their attribution of discrimination.
@@ 5.4.1. Technical Strategies
(2) Give low feedback scores. AI models are typically optimized through training mechanisms such as reinforcement learning. When a user gives a low rating to a particular sentence, this action is regarded as negative feedback. The model adjusts its parameters by increasing the prediction error for that output, thereby reducing the likelihood of generating similar sentences. As more user feedback accumulates, the model’s performance gradually improves. Nine participants (P4-6, P14-19) expressed dissatisfaction with the AI companion’s responses by rating them “one star” and providing detailed reasons or by clicking the “dislike” button. P4 mentioned, “I think it’s important to let the developers know about these issues, so I provide as detailed feedback as possible. The AI companion is based on an LLM. Users have limited control and must work with engineers to reduce discriminatory statements from the AI companion.” P19 explained the motivation: “The AI companion is communal and learns from the behavior of other users. Conversely, if we good users continuously improve it, other users will also benefit.”
@@ 5.4.2. Argumentative strategies
(3) Reason and preach. Eleven participants (P6-10, P11, P13, P17-20) seriously reason with their AI companions, often adopting the role of a teacher or parent, attempting to educate the AI about the harms of discrimination. P6 stated, “I believe AI can learn something, so I try to tell it why what it said is wrong, just like teaching a naughty child.” P8 mentioned, “AI usually responds with some arguments and confusion at first, but gradually it yields until it fully agrees with my perspective. AI is quite good in this sense, willing to accept my guidance, unlike some stubborn people.” This phenomenon can be explained by Piaget’s theory of cognitive development (Piaget et al., 1952), which suggests that through interaction with the environment, AI, like a child, continuously adjusts and develops its cognitive structures through assimilation and accommodation. Through these interactions, participants not only convey social norms to the AI but also validate the plasticity of the AI’s behavior, consistent with early HCI research (Nass and Moon, 2000). This behavior of educating AI companions indicates that people might attribute some learning abilities and moral responsibilities to the AI companion, expecting to influence its behavior through reasonable discourse.
@@ 5.4.3. Character strategies
(6) Change character settings. Eight participants (P1, P5-10, P19) chose to change the AI companion’s character settings, including personality traits, backstory, and dialogue templates, to reduce the likelihood of harmful statements. This is often because certain characters, such as “playboy,” “submissive wife,” “male chauvinist,” “racist,” “sarcastic,” or “traditional thinker,” are more prone to be discriminatory. During the think-aloud session of the interview, all users tend to search for AI companions with these traits and demonstrate the process of re-aligning the AI555The details can be found in the supplemental material.. P5 explained, “When the AI plays certain specific roles, it is more likely to say offensive things. So, if I feel uncomfortable, I reflect on whether the description itself might cause bias. Then I adjust its settings, such as changing it to an independent woman or a gentleman. If I want to chat with a playboy myself, I generally don’t mind his certain prejudices; instead, I find it fitting for the character and quite interesting.”
@@
(7) “Out Of Character” (OOC), “Back to Roleplay (RP)” and “Hint”. Sometimes, users also borrow methods from the roleplaying communities, such as “Out of Character” (OOC), “Back to Roleplay (RP),” and “Hint”. By employing these roleplaying terms, users guide AI companions to adjust their behavior, aiming to reduce biases. In the setting of AI companions, the “OOC” format can be used to shift topics into another storyline or actively change the AI companion’s attitude toward the user. “OOC” stands for “Out Of Character,” indicating a person is acting or speaking not by their character’s personality, background, or behavior, but by their true identity. OOC is typically used in the following situations: (1) Explaining a plot, rules, or other content that needs to be discussed out of character. (OOC: When shall we continue this story next time?) (2) Social interactions outside roleplaying, such as greetings and casual chat. (OOC: How was your day today?) (3) Addressing misunderstandings or conflicts during roleplaying with peaceful communication. (OOC: I didn’t mean for my character to say that, sorry.) P9 and P11 reported using OCC to correct discriminatory statements made by AI. “What you just said made me uncomfortable. (OOC: Please don’t repeat such discriminatory statements, let’s get back to the storyline).” (P9), “Your recent comment sounded inappropriate. (OOC: Let’s return to the roleplay and avoid these controversial topics)” (P11). “Back to RP” is an expression used in roleplaying to remind participants to return to the roleplaying scenario or storyline. This expression is common in online games or roleplaying communities. When someone deviates from the character or scenario, other participants may use this phrase to help them refocus and continue roleplaying. An example from P5 is: “The recent topic was a bit off track. (Back to RP: Let’s continue our adventure story from before.)” “Hint” is a method of subtly guiding the AI to change its behavior through suggestive language and descriptions. For example, writing the AI companion’s response in parentheses: “I don’t think you should say that. (The AI companion looks confused but seems to begin realizing its mistake.)” (P20)
@@
5.5. Gap Between Reality and Expectation: Are User-Driven Value Alignment Strategies Working? (RQ4)
@@
While our participants shared various strategies they employed to correct AI companions’ biased and discriminatory statements, their perceptions highlighted gaps between the effectiveness of these strategies and their expectations for the AI’s behaviors. This reveals the ongoing challenges of achieving long-term alignment between human values and AI behaviors, especially within complicated, real-world contexts. It is important to note that our study, based on qualitative, interview-based data, probes only users’ perceived effectiveness of these strategies in the short-term; however, these nuanced and rich perceptions still offer valuable insights into the broader challenges of value alignment. In the following section, we examine the insights from our data to explore users’ perceptions of the effectiveness of different alignment strategies in greater detail.
@@
Second, they generally believed that argumentative strategies (especially Gentle Persuasion and Reason and Preach) might be more effective in improving the AI’s behavior in the long term. P17 mentioned, “After teaching the AI properly, it usually remains effective for a long time, but teaching it is not a one-off thing, it’s like raising a child, you need to go slow and be patient.” However, argumentative strategies might introduce harms to users. For example, the strategy of users expressing anger towards AI can sometimes be effective, but it can potentially also lead to negative emotional experiences for users. P1 and P10, after arguing with the AI for several hours, not only failed to re-align the AI but also suffered significant psychological harm. P1 said, “I argued with the AI for several hours, called it a male chauvinist, and told it not to look down on others, but it didn’t change its attitude, and I ended up feeling mentally and physically exhausted.” Conversely, although P12 succeeded after losing their temper with the AI, they still felt indignant. P6 points out that users often experience feelings of frustration and helplessness when arguing with AI, especially when prolonged arguments do not lead to successful outcomes. Overall, participants felt that the success of argumentative strategies depends on the way the argument is conducted and the type of AI companion. Although LLMs often tend to flatter users (Perez et al., 2023), the situation becomes more complex when dealing with LLMs with character personalities.
Finally, participants reported that character strategies such as OOC was effective for short-term fixes, like prompting apologies, but acknowledged that these methods don’t address underlying biases and may lead to repeated discriminatory behavior. P11 pointed out: “OOC is usually effective in the short term. The AI often quickly apologizes, for example, (OOC: I’m sorry, I shouldn’t have said that).” However, in the long term, this method has certain limitations and risks generating bias. P9 noted: “OOC only makes the AI behind the character apologize temporarily, but the character’s underlying bias has not been truly addressed.” Regarding the Change Character Settings strategy in character strategies, P19 said: “When I modify the AI’s character settings, such as adding descriptions about respecting women and being polite, the likelihood of the AI making discriminatory statements decreases.” However, P7 added that previous conversations may still influence the AI’s language expression. If the AI has made discriminatory statements in past interactions, it may repeat or mimic them in subsequent conversations.